OpenStudy (amarab):

6. What is the solution? 1/2x^2–x+5=0 possible answers: 1 ± i sqrt 11 –1 ± i sqrt 9 1 ± i sqrt 9 –1 ± i sqrt 11 i would like step by step maybe to help me through this

OpenStudy (anonymous):

i take it this is \[\frac{1}{2}x^2-x+5=0\] if so make your life easier and multiply by 2 first, to get \[x^2-2x+10=0\] then it is much nicer you can either use the quadratic formula, or since \(-2\) is even it is simpler to complete the square

OpenStudy (anonymous):

you would start by subtracting 10 from both sides and write \[x^2-2x=-10\] than complete the square by taking half of 2 which is 1, and write \[(x-1)^2=-10+1=-9\] take the square root, get \[x-1=\pm\sqrt{-9}\] \[x-1=\pm3i\] and finally \[x=1\pm3i\]

OpenStudy (anonymous):

i see in your answer choices you have \(1\pm i\sqrt{9}\) but no one write \(\sqrt{9}\) since it is clearly 3

OpenStudy (amarab):

thank you soooo much! that helped out A LOT